How do I write the equation necessary to find out when my age will be three times that of my grand daughter? Our current ages are 57 and 4. Thank you for your help
Found 2 solutions by solver91311, pedjajov:
Answer by solver91311(24713)



Let represent the number of years between now and the time that you are three times her age. Then is your age at that time and is her age at that time and we want to know the value of when your age is 3 times hers:



Just solve for .

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by pedjajov(51)

If your ages now are 57 and 4 than in x years you(G) and your granddaughter(D) will be:
:
G = 57+x
D = 4+x
:
At the same time, after x years, you should be 3 times older :
:
G=3D
:
57+x=3(4+x), this is equation you're looking for.